ON THE NEW PRECISE SPECTROSCOPIC STUDY OF THE CH_{3}D MOLECULE

O. N. ULENIKOV, \underline{G. A. ONOPENKO}, N. E. TYABAEVA, Tomsk State University, Tomsk, Russia; J. SCHRODERUS, S. ALANKO, and R. ANTTILA, University of Oulu, Finland.

On the basis of recorded (Oulu, Finland) high resolution Fourier transform spectra (ν_{3}, ν_{5}, and ν_{6} bands) and of ”supercombination differences” method^{a} A_{1}–A_{2} (K=3) splittings were first time determined in the ground vibrational state of the ^{12}CH_{3}D molecule for the states with 10 < J < 22. Problem of simultaneous determination of the ϵ̃ and h̃_{3} parameters which are responsible for such splittings is solved on the basis both of theoretical calculation and of analysis of resonance interactions in the recorded experimental spectra. Numerous peculiarities in the spectra of the ν_{3}, ν_{5}, and ν_{6} bands (in particular A_{1}–A_{2} splittings for the K=1,2,3,4,5,6,7, and other effects) are discussed. Analogous problems for the ^{13}CH_{3}D molecule are considered.

aO.N. Ulenikov, S. Alanko, M. Koivusaari, and R. Anttila, Chem. Phys. Lett. 268, p.242-248 (1997)